OpenStudy (anonymous):

2x=1-3y, 3y^2-x^2=2 Simultaneous equations answer+solutions please (NO EXPLANATION NEEDED)

OpenStudy (tyteen4a03):

If you're only looking for answers and solutions, you're looking for Wolfram|Alpha, not OpenStudy. OpenStudy is not a place for you to skip your homework. Notice the similarity in these two equations - they both have the term 3y. We need to do some work to make 3y usable (notice the negative sign attached to it), so move the equation around: \(3y = 1 - 2x\). Now, we can substitute this to the circle equation \((1 - 2x)^2-x^2=2\). Now solve for x like how you solve quadratic equations, then use the newfound value of x to find y using the line equation. Notice that there are two answers - make sure you solve for y for both roots of \((1 - 2x)^2-x^2=2\).

OpenStudy (anonymous):

Well bro, it's 12:16 AM, I'm not even half done with my homework and wolfram alpha shows no solutions. Sorry bro if it ticks you off, but I'm pretty winded-out man.

OpenStudy (tyteen4a03):

@Matt71 Wolfram|Alpha does give out solutions to this set of equations. http://www.wolframalpha.com/input/?i=2x%3D1-3y%2C%203y%5E2-x%5E2%3D2&t=crmtb01&f=rc I've already solved half of the equation for you, you just need to quickly look over the information and solve the rest of it, and you won't have to rely on OpenStudy anymore. Also, it is also 12AM here. ;)

OpenStudy (anonymous):

Dang, thanks man